# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

### v1

@@ 2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the interception $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$. Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant. Now, for any $p=x\cdot\Big{(}{\rm SO}(2)\times{\rm SO}(3)\Big{)}\in G_{5,2}$ define $(\omega,\real\Omega)$ by
@@
where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$. ∎
The following result is mentioned in [kerr1998new]*§4, for sake of completeness, we include the detail proof of it.
@@
Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphism on ${\mathfrak{m}}$ given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$. Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
A straightforward computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
@@
The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ are described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
@@
Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written by
@@
From (2.6) is easy to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on ${\mathfrak{m}}$ is given by
@@
The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure, if the induced symmetric bilinear form
$B_{ij}=\Big{(}(e_{i}\lrcorner\varphi)\wedge(e_{j}\lrcorner\varphi)\wedge\varphi\Big{)}(e_{1},...,e_{7})$
is positive (or negative) defined. Explicitly, we obtain that for $k=2,3,4$
$B_{11}=6\widetilde{z}^{3},\quad B_{kk}=6\widetilde{z}(\widetilde{b}^{2}+\widetilde{a}\widetilde{x}),\quad B_{k+3k+3}=6\widetilde{z}(\widetilde{a}^{2}+\widetilde{b}\widetilde{x}),\quad B_{kk+3}=3\widetilde{z}(\widetilde{a}\widetilde{b}-\widetilde{x}\widetilde{y})$
@@
$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{missing}}{det}\nolimits(B)^{\frac{1}{9}}},$
hence, by comparing $g$ with (2.7), we obtain the condition in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
@@
From Lemma (2.3), the following result describe all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
@@
From a straightforward computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed. Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14). ∎
@@
The $S^{1}$-family given in (2.14) correspond to the general construction given [alexandrov2012]*(4.28).
@@ 2.2 Homogeneous Sasakian metrics on $V^{5,2}$
In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$. Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metrics structures, cf. Section 1. Secondly, in Theorem 2.2, we provide the constraints in the metric to get a Sasakian structures. And finally, we relate the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric.
@@
First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16). Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3)$-invariant endomorphism $\Phi$ on ${\mathfrak{m}}$ such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$. From (2.9) we have
@@
Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $-\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
@@
For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $[\Phi Y,Y]$ and $[X,\Phi Y]$ are generated by $\xi$ (cf. (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$. So, (2.21) holds by noting that
@@
In this case, the $2$-parameter family of coclosed $\mathrm{G}_{2}$-structures is
@@
To proof item $(1)$, we compare (2.13) and (2.22), thus
@@ 3 $\mathrm{G}_{2}$ and self dual contact instantons on $V^{5,2}$
Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution $H=\ker\eta\subset TM=H\oplus\xi$. Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$. Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms. We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes by (cf. [portilla2023, Section. 2.1])
@@
With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in (2.11) we have that: If $A$ is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
@@
We know that in local coordinates, a basis for self dual contact $2$–forms is given by (cf. [portilla2023, Section 2])
@@
For $\mathop{\mathrm{Re}}\Omega_{0}$ and $\mathop{\mathrm{Im}}\Omega_{0}$ given in Lemma 2.1, one verifies that $w_{i}\wedge\real\Omega_{0}=w_{i}\wedge\mathop{\mathrm{Im}}\Omega_{0}=0$ for each $i=1,\cdots,8$ in (3.3). Hence, if $A$ is a SDCI, then its curvature $F_{A}$ is locally generated by $w_{i}$’s in (3.3). Thus, for the $\mathrm{G}_{2}$-structure (cf. Corollary 2.8):
@@
where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$. On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
@@
On the other hand, we have the $\mathrm{G}_{2}$-invariant decomposition $\Omega^{2}(TM)=\Omega^{2}_{7}\oplus\Omega^{2}_{14},$ where $\Omega^{2}_{14}\simeq{\mathfrak{g}}_{2}$. Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$. It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$, hence the same is true for $\Omega^{2}_{6}$.
@@
Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], with the Sasakian structure $(\eta,\xi,g,\Phi)$ given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection. If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton then $A$ satisfies the equation (1.2), i.e. it is a $\lambda$-contact instanton .
@@
Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are solutions to the contact instanton equation. ∎
@@
We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU$(n)$-structure. These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
@@ 4 Homogeneous principal bundles on $V^{5,2}$
First, we recall some notation and results about homogeneous principal bundle, following [cap2009]*Section 1.4. Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, $L\colon G\times M\to M$ the canonical left action and $p\colon G\to M$ an $H$–principal bundle given by the canonical projection.
@@
A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left $G$–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
@@
[cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup. Let $P\to G/H$ a homogeneous $K$-principal bundle, then there exist a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$. The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$. The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exist $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant, if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.

### v2

@@ 2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the intersection $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$. Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant. Now, for any $p=x\cdot\Big({\rm SO}(2)\times{\rm SO}(3)\Big)\in G_{5,2}$ define $(\omega,\real\Omega)$ by
@@
where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ . Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$. ∎
The following result is mentioned in [kerr1998new]*§4. For the sake of completeness, we include the detailed proof.
@@
Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphisms on $\mathfrak{m}$} given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$. Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
A direct computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
@@
The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ is described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
@@
Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written as
@@
From (2.6) it is straightforward to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on $\mathfrak{m}$} is given by
@@
The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure if the induced symmetric bilinear form
$B_{ij}=\Big((e_{i}\lrcorner\widetilde{\varphi})\wedge(e_{j}\lrcorner\widetilde{\varphi})\wedge\widetilde{\varphi}\Big)(e_{1},...,e_{7})$
is positive (or negative) definite. Explicitly, we obtain that $B_{11}=6\widetilde{z}^{3}$ and for $k=2,3,4$
[…]
@@
$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{det}}\nolimits(B)^{\frac{1}{9}}},$
hence, by comparing $g$ with (2.7), we obtain the conditions in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
@@
From Lemma (2.3), the following result describes all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
@@
From a direct computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed. Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14). ∎
@@
The $S^{1}$-family given in (2.14) corresponds to the general construction given in [alexandrov2012]*(4.28).
@@ 2.2 Homogeneous Sasakian metrics on $V^{5,2}$
In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$. Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metric structures, cf. Section 1. Secondly, in Theorem 2.2, we provide the constraints in the metric to get a Sasakian structures. And finally, we relate the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric.
@@
First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16). Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3))$-invariant endomorphism $\Phi$ on $\mathfrak{m}$} such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$. From (2.9) we have
@@
Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
@@
For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $\Phi Y,Y]$ and $X,\Phi Y]$ are generated by $\xi$ (cf. (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$. So, (2.21) holds by noting that
@@
In this case, the 22-parameter family of coclosed $\mathrm{G}_{2}$-structures is
@@
We compare (2.13) and (2.22), thus
@@ 3 $\mathrm{G}_{2}$ and self dual contact instantons on $V^{5,2}$
Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution given by $H=\ker\eta\subset TM=H\oplus\xi$. Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$. Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms.
We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes as (cf. [portilla2023, Section. 2.1])
@@
With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in Corollary 2.8 we have that: If AA is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
@@
We know that in local coordinates, a basis for self dual contact 22–forms is given by (cf. [portilla2023, Section 2])
@@
For $\mathop{\mathrm{Re}}\Omega_{0}$ and $\mathop{\mathrm{Im}}\Omega_{0}$ given in Lemma 2.1, one verifies that $w_{i}\wedge\real\Omega_{0}=w_{i}\wedge\mathop{\mathrm{Im}}\Omega_{0}=0$ for each $i=1,\cdots,8$ in (3.3). Hence, if $A$ is a SDCI, then its curvature $F_{A}$ is locally generated by $w_{i}$’s in (3.3). Thus, for the $\mathrm{G}_{2}$-structure (cf. Corollary 2.8)
@@
where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$. On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into the $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
@@
On the other hand, we have the $\mathrm{G}_{2}$-invariant decomposition $\Omega^{2}(TM)=\Omega^{2}_{7}\oplus\Omega^{2}_{14},$ where $\Omega^{2}_{14}\simeq{\mathfrak{g}}_{2}$. Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$. It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$. Hence, the same is true for $\Omega^{2}_{6}$.
@@
Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], the Sasakian structure $(\eta,\xi,g,\Phi)$ is given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection. If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton, then $A$ satisfies the equation (1.2), i.e. it is a $\lambda$-contact instanton .
@@
Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are also solutions to the contact instanton equation. ∎
@@
We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU(n)(n)-structure. These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
@@ 4 Homogeneous Principal Bundles on $V^{5,2}$
First, we recall some notation and results about homogeneous principal bundles, following [cap2009]*Section 1.4. Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, let $L\colon G\times M\to M$ be the canonical left action and $p\colon G\to M$ an HH–principal bundle given by the canonical projection.
@@
A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left GG–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
@@
[cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup. Let $P\to G/H$ be a homogeneous $K$-principal bundle, then there exists a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$. The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$. The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exists $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.